OpenStudy (anonymous):

which equation is equivalent to square root of 24 plus x equals twelve? x squared + 24x + 168 x squared - 24x + 120 x squared - 24x + 168

OpenStudy (anonymous):

is it: \[\sqrt(24) + x = 12?\] or \[\sqrt(24+x) = 12?\]

OpenStudy (anonymous):

sstarica - I believe it is the former. Otherwise you wont get the X^2 term in the result

OpenStudy (anonymous):

hmm, alright then if so then the answer is B) the second part : Solve for x ^_^: \[\sqrt(24) + x = 12 \rightarrow 24 + x^2= 144\] hmm there's something wrong, is it 24x? or 24?

OpenStudy (anonymous):

24x :) When you square √24)+x=12 you get \[(√24+x)^2 = 12^2\] that will give you a 24x

OpenStudy (anonymous):

are you sure? that would give 48x lol >_< okay, it's either I'm going off, or this question is missing a part lol